EVA Report – April 13th

Crew 245 EVA #2 Report 13-04-2022
SOL: 3
EVA #: 2
Author: Simone Paternostro

Purpose of EVA: We performed another EVA to collect geological and biological samples in three locations in the vicinity of the Habitat and East area (Pooh’s corner) and continue to collect data with wearable sensors on arms and shoes.

Start time: 10:00
End time: 12:30

Narrative: The crew collected the first samples in the area next to the solar panels on the east side. After that, they took two rovers (Spirit and Opportunity) and reached two additional points on the East side of the Habitat (one near the Marble ritual and one in Pooh’s corner)

Destination: Marble ritual and Pooh’s corner

Coordinates (UTM NAD27 CONUS): 4250946.023 518266.204
4251040.273 518584.714
4250991.861 518827.441

Participants: Simone Paternostro (EVA Commander), Nadia Maarouf (EVA HSO), Benjamin Pothier
Road(s) and routes per MDRS Map: Cow Dung road and Lowell Highway

Mode of travel: Walking and Rover

Journalist Report – April 13th

"Getting grounded" Journalist Report
Crew 245 Journalist report 13 April 2022
Sol: 03
Name of person filing the report: Dr. Benjamin Pothier

This is Sol 3 for our team and we have already reached our cruise speed in terms of carrying out and planning our experiments in the field. Despite the constraints of the Martian environment and the food and water restrictions, the crew has adapted well to our new normal. I personally plan to take my next shower in a couple of days and we all understand that every resource is precious here. It’s all about ISRU, or In-Situ Resource Utilization and I’m sure that we will all be more and more creative in the next couple of days with what we have at hand in order to enhance and optimize our performances, be it during cooking, DNA analysis or drone photogrammetry.

Everyone is important in our team because the success of our mission, from our comfort to plain survival, depends on each and every skill. If Buckminster Fuller spoke about the "Spaceship Earth" in 1969, our station is our Spaceship, a very grounded one, because the environment around us is hostile, and in Space, everything tries to kill you. We don’t go outside without our spacesuits and we need to plan and rationalize our use of each and every resource that is available to us.

I joined our Crew Scientist Simone and our Health and Safety officer Nadia during an Extra-Vehicular Activity this morning, documenting their process of soil sampling for an upcoming DNA analysis experiment. This afternoon our Commander Vittorio with our Crew Engineer Luca and our Executive Officer Luca installed a ground station antenna that should let us communicate with some satellites in the next couple of days.

I can’t wait to collaborate further with all those crew members on the several projects that we have planned. Civilization was always about collaboration on Earth, and it is the same in our small resilient community. We are slowly getting grounded on Mars and I can’t wait for the next sunrise.

Crew Journalist Over and out.

EVA Report – April 13th

Crew 245 EVA Report 13-04-2022
SOL: 3
EVA #: 3
Author: Vittorio Netti

Purpose of EVA: Install the antenna ground component

Start time: 14:30
End time: 17:00

Narrative: The purpose of this EVA was to install the ground component of the satellite antenna for the D-Orbit experiment. The tasks included the positioning on the ground of a 5.5 feet metal pole and the deployment of three metal cables connected to three concrete counterweights to stabilize the Antenna in case of high winds. The manual work was particularly challenging due to the need to tighten 36 metal M2 bolts on the cables.

We also filled two plastic bags with Martian regolith to provide additional stabilization to the feet of the pole. At the end of the EVA, we dedicated some time to looking for a lost suit sensor from the previous EVA crew, unsuccessfully.

Destination: Outside of the RAM

Coordinates (UTM NAD27 CONUS): 4250929.883 N; 518164.975 E

Participants: Paolo Guardabasso (EVA Commander), Vittorio Netti (EVA HSO), Luca Rossettini

Road(s) and routes per MDRS Map: outside of the RAM

Mode of Travel: Walking

Commander out.

Sol Summary – April 13th

Crew 245 Sol Summary Report 13-04-2022
SOL: 3
Summary Title: Busy day
Author’s name: Paolo Guardabasso

Mission status: All experiments are nominal and running.

Sol Activity Summary: Busy day for the crew, between EVAs and the last breaking of Sim to retrieve the last heavy package (Industrial 3D Printer).

Look Ahead Plan: Tomorrow a morning EVA will be performed to collect some more samples, and the afternoon will be dedicated to experiments inside the station.

Anomalies in work: A sensor was lost in the field today, it’ll be recovered tomorrow. If it won’t rain, it might survive the night. Otherwise, we’ll do a short ceremony and a minute of silence for this sacrifice to science.

Weather: Very pleasant in the morning, getting windy in the afternoon. Temperature -1 (low) and 12 (high), Wind speed 35 (avg) and 45 (max).

Crew Physical Status: All nominal.

EVA: Today 2 EVAs were performed: a morning EVA for sampling and an afternoon EVA to fix and assemble the base for our Ground Station antenna.

Reports to be filled:
– HSO report
– Operation report
– Journalist
– Picture of the day
– Some pictures
– EVA request
– EVA report (2)

Support Requested: None

Operations Report – April 13th

Crew 245 Operations Report 13-04-2022
SOL: 3
Name of the person filing report: Luca Rossettini

Non-nominal systems:
Heating system.
Suit 8 battery.
One suit charger.

Notes on non-nominal systems:
The heating system is not working. We are using the heater from the RAM at the upper deck.
Suit 8 battery has been substituted.
Faulty suit charger: a label was put signaling that is not working.


Spirit rover used: Yes
Hours: 167.7
Beginning charge: 100%
Ending charge: 96%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 93.9
Beginning charge: 100%
Ending charge: 97%
Currently charging: Yes

Curiosity rover used: No
Hours: 182.7
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Perseverance rover used: No
Hours: 235.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

General notes and comments: None

Summary of Hab operations: –
Water (static tank): 415 gallons
Water (loft tank): 12 gallons
Water Meter: 157522,9
Static to Loft Pump used: Yes
Static tank pipe heater: Off
Static tank heater: Off
Toilet tank emptied: No

Summary of internet: nothing to report

Summary of suits and radios:
The battery of suit 8 has been found faulty and has been substituted.
One suit charger has been found broken. We put a label on it.

Summary of GreenHab operations:
WATER USE: 0 gallons
Heater: Off
Supplemental light: Off

Summary of ScienceDome operations: Nothing to report
Dual split (Heat/AC): Off

Summary of RAM operations: We prepared the systems for the satellite experiment.

Summary of any observatory issues: Not used, nothing to report

Summary of health and safety issues:

Questions, concerns and requests to Mission Support: –

Supplemental Operations Report – April 13th

Supplemental Operations Report 13 April 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Hab furnace

Action taken for non-nominal systems: Ordered a new thermostat to see if that is the problem as the heater just appears dead. If that does not fix the problem, I will just be leaving it as we will be installing a new furnace for next season.

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 59 %

Propane Reading, director tank— 53%

Propane Reading, intern tank— 69 %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 350 gallons

Sojourner rover used: no

Hours: 167.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes
Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? Yes, to town

Luna used and why, where? No

Campus wide inspection, if action taken, what and why? The Guard rail on the front porch was damaged in the windstorm. The crew repaired it.

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of Outpost operations: New toilet arrived early. Hoping to install it tomorrow!

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site